OpenStudy (jozelynw):

Iris has been studying an invasive population of snails. This particular snail has no local predators, so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. 2.) A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population.

OpenStudy (jozelynw):

@pooja195 please help me

OpenStudy (anonymous):

\[g(x)=200(1.03)^{2x}=200(1.03^2)^x\] Growth rate formula \[g(x)=A(1+r)^x\] A = initial amount, r = growth rate

OpenStudy (jozelynw):

the answer is g(x)=200(1.0609)^x. I need help with the second part of the question @peachpi

OpenStudy (jozelynw):

@dan815

OpenStudy (jozelynw):

@readergirl12

OpenStudy (jozelynw):

@avis

OpenStudy (anonymous):